CASES99: Logbook Entries

CASES99: Site tower Messages, 75 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
193 Tue 16-Nov-1999final configuration for all ADAMStowerknudson
189 Fri 29-Oct-1999moving 10m boomtowerlundquist
187 Thu 28-Oct-1999Licor zerotowerdelany
185 Thu 28-Oct-1999New winch cabletowerdelany
180 Tue 26-Oct-1999Ragwort ventilationtowerdelany
179 Tue 26-Oct-1999Hauling cabletowerdelany
176 Tue 26-Oct-1999Ventilation of Ragworttowerdelany
172 Mon 25-Oct-1999Wind directions on towertowerdelany
170 Sun 24-Oct-1999Zeroed Licortowerdelany
167 Sun 24-Oct-1999Link to tower down / fiber damagetowersemmer
166 Sat 23-Oct-1999All ADAMS reset at 16:09 localtowersemmer
154 Fri 22-Oct-1999Floppy used spacer used on ragwort lidtowersemmer
153 Thu 21-Oct-1999Comb spacer for Ragwort changedtowerdelany
152 Thu 21-Oct-1999Orientation of sensor boomstowerdelany
151 Thu 21-Oct-1999comb orientation at ragwort changedtowersemmer
148 Wed 20-Oct-1999Cleaned 5m KH2O at tower tower
147 Wed 20-Oct-199915m TRH replacedtowersemmer
141 Tue 19-Oct-1999Mounted new h2o/co2 sensor at 20mtowerhorst
138 Mon 18-Oct-1999daisy crashed and rebootedtowerhorst
137 Mon 18-Oct-1999Removed 20m h2o/co2 sensortowerhorst
136 Mon 18-Oct-1999Cleaned radiometers at 50m and PAM stationstowerhorst
130 Sun 17-Oct-199920m hot wire installed this afternoon.towerhorst
126 Sun 17-Oct-1999Krypton back on the air at 20 meterstowersemmer
125 Sun 17-Oct-1999CO2 back on at 20 meters towersemmer
121 Sun 17-Oct-1999Krpyton at 20m off the air at ~ 15:30 local timetowersemmer
120 Sun 17-Oct-1999work on CO2 sensortowersemmer
118 Sun 17-Oct-199920m sonic noisy on v axistowerhorst
115 Sun 17-Oct-1999h2o.k.20m back uptowersemmer
114 Sun 17-Oct-1999h2o.k.20m not workingtowersemmer
113 Sat 16-Oct-1999Tests on cosmos analog channel 123towersemmer
112 Sat 16-Oct-1999Hotfilm sheath signal is backtowersemmer
111 Sat 16-Oct-1999Signals from Hotfilm sheath down for awhiletowersemmer
110 Sat 16-Oct-1999Ragwort back uptowersemmer
108 Sat 16-Oct-1999raising/lowering 40m boomtowerlundquist
107 Sat 16-Oct-1999Cycled power on 10m sonic on 55m towertowerhorst
106 Sat 16-Oct-199920m co2 statustowersemmer
105 Sat 16-Oct-199940m co2 power supply taken downtowersemmer
104 Sat 16-Oct-1999Ragwort off the airtowersemmer
103 Fri 15-Oct-1999moving 40m boomtowerlundquist
101 Fri 15-Oct-1999Cycled power on 10m sonic on 60m towertowerhorst
98 Thu 14-Oct-199920m sonic died; replaced w/ATI (and KH2O from Station 6)towerhorst
97 Thu 14-Oct-1999Raising 10m and 40m booms to remove hotwirestowerlundquist
95 Thu 14-Oct-1999No licor cal gasestoweroncley
94 Wed 13-Oct-1999hotwire replacement/look on 55-m tower (Lundquist)towerlemone
92 Wed 13-Oct-1999tour up 55m towertoweroncley
89 Tue 12-Oct-1999Replaced hygrothermometer fan at 15mtowerhorst
85 Tue 12-Oct-1999towermiller
82 Tue 12-Oct-1999ati.55m power cycletoweroncley
81 Mon 11-Oct-1999new Ironics card at ragworttowermaclean
80 Mon 11-Oct-1999ragwort overheatingtowermaclean
78 Mon 11-Oct-199920m H2O/CO2 reinstalledtoweroncley
77 Mon 11-Oct-199920m boom up/CO2 removedtoweroncley
75 Mon 11-Oct-1999Cleaned 5m KH2Otowerhorst
73 Sun 10-Oct-199920m ATI was at 10 samp/stoweroncley
72 Sun 10-Oct-1999replaced 10m ATItoweroncley
70 Sun 10-Oct-199920m boom orientationtowerhorst
67 Sat 09-Oct-19995m tower down timestowermiller
62 Sat 09-Oct-1999raised and lowered 20m boomtowerlundquist
60 Sat 09-Oct-1999swapped prop at 15mtoweroncley
59 Fri 08-Oct-1999co2.20m fixed itself?toweroncley
58 Fri 08-Oct-1999Moving the 40 m boomtowerlundquist
57 Fri 08-Oct-1999new krypton at 55mtoweroncley
54 Fri 08-Oct-1999approximate licor calibrationtoweroncley
52 Fri 08-Oct-1999Licor co2 installedtoweroncley
51 Thu 07-Oct-1999daisy problemstowermaclean
48 Tue 05-Oct-1999co2 turned offtoweroncley
44 Mon 04-Oct-1999Movement of boomstowerhorst
39 Thu 30-Sep-1999ADAM statustowermaclean
38 Wed 29-Sep-1999ADAM statustowermaclean
32 Tue 28-Sep-1999ADAM statustowermaclean
29 Sun 26-Sep-1999ADAM statustowermaclean
28 Sat 25-Sep-1999sonic parameterstoweroncley
27 Sat 25-Sep-1999ADAM status, turned off chem shelter ACtowermaclean
24 Sat 25-Sep-1999ADAM statustowermaclean
22 Fri 24-Sep-199955m tower guy wire tensionstowerknudson


193: ADAMS, Site tower, Tue 16-Nov-1999 11:11:05 MST, final configuration for all ADAMS
final resting place for all ADAM cards at teardown


   Ragwort Matrix card s/n 9006
   16 channel Ironics serial card s/n 821003
   Datel Dvme 601 A/D card s/n 320689
   signal conditioner card # 2
   NCR power supply s/n 000371
   chassis # 1  (missing insulation on enclosure lid)
   Teca cooler # 1 s/n 9121083  NSF 23024


   Daisy Matrix card s/n 9010
   8 channel Ironics serial card (chicory) s/n 014029
   Datel Dvme 601 A/D card s/n 440690
   NCR power supply (belongs to PSU) s/n 002672
   chassis # 4 (missing insulation on enclosure lid)
   Teca coller # 4 s/n 8940208


   Cosmos Matrix card s/n 940211101
   16 channel Ironics serial card s/n 129043
   Datel Dvme 601 A/D card s/n 400190
   Datel Dvme 641 expansion card s/n 001231288
   signal conditioning cards # 3, # 7, and a no number board (belongs to PSU)
   NCR power supply s/n 000307
   chassis - no number (belongs to PSU)
   Teca cooler s/n 9421060 (belongs to PSU)


   Marigold Matrix card s/n 9009
   8 channel Ironics serial card (Edelweiss) s/n 67356
   Datel Dvme 601 A/D card 320589
   NCR power supply (marked dead) s/n 000298
   chassis # 2
   Teca cooler # 3


   Fester Matrix card s/n 9013
   16 channel Ironics serial card s/n 821022
   Datel Dvme 601 A/D card s/n 000940694
   signal conditioning card # 6
   NCR power supply s/n 000238
   chassis # 3
   Teca cooler # 2


189: Hotwire, Site tower, Fri 29-Oct-1999 11:33:40 CDT, moving 10m boom
From 10:50-11:10, the 10m boom was raised so that I could remove the hotwire
(#69) at 10m. I had intended to go up to 40m as well, but winds were too
strong.
187: CO2, Site tower, Thu 28-Oct-1999 12:30:37 CDT, Licor zero
At ~ 1200 the Licor was zeroed for ~ 15 minutes
185: Tower, Site tower, Thu 28-Oct-1999 12:26:54 CDT, New winch cable
Between 0930 and 1030 Tony, Cathy, Sean and Julie strung a new winch cable for 
the main tower. 
180: ADAMS, Site tower, Tue 26-Oct-1999 20:22:09 CDT, Ragwort ventilation
Inspection at ~1500, 26 Oct''99, indicated that the disc inserted in the morning
to hold open the lid of Ragwort had been placed so that it held the radiation 
shield of the ADAM rather than the lid. This was fortuatous as this spacing 
resulted in good daytime ventilation (max Tbox ~ 32 C for a hot day ). This
arrangement will be maintained for nighttime ventilation
179: Tower, Site tower, Tue 26-Oct-1999 20:15:11 CDT, Hauling cable
Monday, 25 Oct''99, in the afternoon, the thermal imaging equipment was winched
to the 58m level of the tower. During the process the hauling cable snagged in
the winch gears and got chewed up. The wire ran through the boom pulleys and 
fell to the ground wiping out the lowest thermocouple but causing no other 
damage.
On Tuesday Tony and Sean inspected the boom and devised the means to thread a 
new cable.
176: ADAMS, Site tower, Tue 26-Oct-1999 13:13:24 CDT, Ventilation of Ragwort
Throughout the nighttime the comb held the lid of Ragwort open ~ 1 cm. 
The temperature only fell to 20C. 
At ~ 0900 the comb was removed and the disc was substituted holding the lid
open ~ 5 cm.
172: PropVanes, Site tower, Mon 25-Oct-1999 10:18:48 CDT, Wind directions on tower
Inspection of the wind direction plots indicate that the wind direction does not change monotonically with height for the tower.

170: CO2, Site tower, Sun 24-Oct-1999 13:01:17 CDT, Zeroed Licor
Between 1230 and 1300 a zero canister was attached to the intake of the 
Licor CO2 sensor.

Front panel display for zero    = - 0.2 
Front panel display for ambient = + 2.4 to 2.5

The pump was running well but I couldn't check the flow as the flow guage
was missing.
Both cal cylinders were empty.
167: Network, Site tower, Sun 24-Oct-1999 09:45:49 CDT, Link to tower down / fiber damage
The link to the tower went down at 3:15 am local.
The problem was a bad fiber in the 300m cable. The
bad fiber is orange. All stations were back on the
air at 9:50 am local.
The 300m cable was checked for cuts. Nicks were
found at 2 locations.
We are now down to 4 fibers with 1 week to go.

166: ADAMS, Site tower, Sat 23-Oct-1999 17:04:27 CDT, All ADAMS reset at 16:09 local
All the ADAMS were reset at 16:09 local.
At this time, I do not know why. They came back
up without the need for human input.


154: ADAMS, Site tower, Fri 22-Oct-1999 10:15:22 CDT, Floppy used spacer used on ragwort lid
The ragwort lid spacer, the comb, was repalced with a
floppy disk at ~10:00 local.


153: ADAMS, Site tower, Thu 21-Oct-1999 17:40:46 CDT, Comb spacer for Ragwort changed
At 1730 the comb spacer for Ragwort was changed from vertical to horizontal.
Note that even when vertical the spacing was not large because plastic comb
was deformed until nearly bent double.
152: Tower, Site tower, Thu 21-Oct-1999 16:36:22 CDT, Orientation of sensor booms
Steve Semmer and Tony Delany shot the angles of the ten east-side booms
between 1500 and 1630 CD Time, Thurs, 21 Oct. The tripod was set up on
the road. The 5m tower sonic anemometers could not be accurately viewed
from that distance.

    55m  Sonic  275 54' 20"     
    50m  Sonic  276 12' 50"
    45m  Prop   276 25' 30"
    40m  Sonic  276 24' 10"
    35m  Prop   277 11' 55"
    30m  Sonic  277 12' 20"
    25m  Prop   276 21' 15"
    20m  Sonic  275 48' 50"
    15m  Prop   277 12' 25"
    10m  Sonic  277 12' 30"

10/23 Cathy and Steve shot the angles of the 2 sonics on the 5 metre tower.
The bottom sonic (.5m) seemed to be unlevel as well as at an angle to the boom.

     5m CSAT	276 42' 30"
    .5m CSAT    278 24' 05"


151: ADAMS, Site tower, Thu 21-Oct-1999 11:53:21 CDT, comb orientation at ragwort changed
The high today is predicted to be high 70s. So Cathy
and I changed the comb orientation from horizontal
to vertical.

148: Kryptons, Site tower, Wed 20-Oct-1999 14:56:16 CDT, Cleaned 5m KH2O at tower
I cleaned the windows of the KH2O at 5m at the main tower, using
a q-tip and distilled water.  I intended to also clean those on the
PAM stations, but realized that the short ladder was not adequate
(duh!).  Since I was too lazy to shlep the large ladder around to
all the PAM stations, I bagged the whole thing.

147: TRH, Site tower, Wed 20-Oct-1999 11:37:37 CDT, 15m TRH replaced
The 15m TRH humidity level has been low compared to the
other sensors on the tower, ~ 2%. The sensor was replaced
with one of the spares. This occurred at ~11:20 local time.
Will monitor its performance.

1500 CDT, Oct 20
This appears to have increased both the T and RH values slightly
and they now both fit into the 55m tower profile better.  Steve
found that the previous 50Y was an older type and the PRT had
been replaced with a smaller, faster-response transducer (during
CASES97?) -- twh


141: CO2, Site tower, Tue 19-Oct-1999 19:07:45 CDT, Mounted new h2o/co2 sensor at 20m
ABLE (Argonne) provided a replacement for the bad h2o/co2 sensor
at 20m.  We mounted this on the tower this evening, raising the 20m
boom around 1710 CDT and lowering it at 1755 CDT.

At this time we removed the KH2O sensor for remounting on Station 6.



138: ADAMS, Site tower, Mon 18-Oct-1999 18:01:18 CDT, daisy crashed and rebooted
Daisy stopped around 16:10 CDT, Oct 18.  We could not 'ping' daisy,
so Steve recycled power at 17:59 CDT.  It was around 16:10 that
we were working on the tower, but it is not apparent how we could
have caused this crash.  We lowered the 20m boom at 16:05 and then
went up to 50m to move the radiometers to the end of their boom.
137: CO2, Site tower, Mon 18-Oct-1999 16:40:30 CDT, Removed 20m h2o/co2 sensor
We have given up on fixing the ANL h2o/co2 sensor on the 20m boom.
We raised the boom at 15:20 to remove it and lowered the boom at
16:05 (after adding another collar to hold the boom pivot in 
position).
136: Radiation, Site tower, Mon 18-Oct-1999 16:38:11 CDT, Cleaned radiometers at 50m and PAM stations
We retracted the 50m radiometers and cleaned them, beginning at 1500 CDT,
Oct 18.  They were cleaned with "de-mineralized" water and kimwipes.

The radiometers were returned to their outboard position at 1615 CDT
(after unfouling the hauling line with the loooon green hook).

The domes of the surface radiation sensors were also cleaned after 
finishing the tower radiometers.


130: Hotwire, Site tower, Sun 17-Oct-1999 19:39:08 CDT, 20m hot wire installed this afternoon.
17 Oct
The 20m boom was raised this afternoon, Oct 17, at 13:13:30 CDT.  
It was lowered at 13:23 CDT.  At this time the hot wire was installed.



126: Kryptons, Site tower, Sun 17-Oct-1999 19:19:37 CDT, Krypton back on the air at 20 meters
The krypton at 20 meters is back on the air. It is the
primary h2o sensor. The CO2 was also remounted and is
on the air. The boom was lowered at 18:25 CDT.




125: CO2, Site tower, Sun 17-Oct-1999 19:17:25 CDT, CO2 back on at 20 meters
The CO2 sensor was remounted at 20 meters after a preliminary
test with the new 5 volt supply. We left the krypton there
as the primary h2o sensor. The boom was lowered at 18:25 CDT.



121: Kryptons, Site tower, Sun 17-Oct-1999 14:49:42 CDT, Krpyton at 20m off the air at ~ 15:30 local time

120: CO2, Site tower, Sun 17-Oct-1999 14:46:38 CDT, work on CO2 sensor
It appears that the problem with the CO2/H2O sensor was
in the 5 volts supply. I modified their power supply so
that we could use a different 5 volt source. The sensor
was deployed for testing at the 20 meter height. After
we collect some data we will decide if it is good or bad.
The sensor was deployed at ~13:50 local time.

118: Sonics, Site tower, Sun 17-Oct-1999 13:03:14 CDT, 20m sonic noisy on v axis
The v axis on the 20m ATIK has been spiking.  We plan to move this unit
back to station 6 this afternoon and mount the sonic removed from 
10 m on Oct 10.  SPO found that this sonic was improperly programmed
as though it were a UW array.  He fixed this and we will mount it at
20m.

115: Kryptons, Site tower, Sun 17-Oct-1999 10:01:46 CDT, h2o.k.20m back up
The krypton at 20m is back on the air. The analog
signal cable was plugged into the wrong channel.

114: Kryptons, Site tower, Sun 17-Oct-1999 09:35:09 CDT, h2o.k.20m not working
The 20m krypton was not working throughout the night.
I may have incorrectly connected it back up after the
breakout box test.

113: ADAMS, Site tower, Sat 16-Oct-1999 16:13:04 CDT, Tests on cosmos analog channel 123
During the day, tests were run to determine the source of noise
on analog channel 123, h2o.k.20m, on cosmos. Below is the list of
tests and the outcome.

(1) switched h2o.5m with h2o.k.20m. So h2o.5m signal was going
into channel 123 and h2o.k.20m was going into 108.
result: noise problem stayed with channel 123.

(2) changed analog/filter board
result: noise on 123

(3) removed special control board near power supply
result: noise on 123

(4) reversed h2o.5m and h2o.k.20m. switched analog breakout boxes
result: noise still on channel 123

112: Hotfilm, Site tower, Sat 16-Oct-1999 16:12:28 CDT, Hotfilm sheath signal is back

111: Hotfilm, Site tower, Sat 16-Oct-1999 15:02:18 CDT, Signals from Hotfilm sheath down for awhile
16 Oct
The signals from the hotfilm sheath are down. They will be
down for a test of an analog breakout box.



110: ADAMS, Site tower, Sat 16-Oct-1999 15:01:25 CDT, Ragwort back up
Ragwort is back on the air. Tom and I decided to leave it
on tonight. There is a slight chance of rain.

108: Hotwire, Site tower, Sat 16-Oct-1999 09:25:32 CDT, raising/lowering 40m boom
At 08:47 CDT, I raised the 40m boom to uninstall the #70 hot wire from the
intercomparison experiment last night with HRDL. By 08:53, it was back 
down. 
107: Sonics, Site tower, Sat 16-Oct-1999 09:20:04 CDT, Cycled power on 10m sonic on 55m tower
We turned off cosmos temporarily at 1350 CDT, Oct 15, for a test of
the filtering on the 20m KH2O.  This stopped the 10m sonic again.
Power was cycled on the 10m sonic at 1715 CDT to restart it.
106: CO2, Site tower, Sat 16-Oct-1999 08:46:59 CDT, 20m co2 status
I ran some tests on the 20m co2 sensor. Based on feedback from
Dave Cook at Argone, it appears the sensor is bad. At one time,
it appeared the 5 volt supply was the problem; however, tests
with the other co2 supply showed this not to be the case. We will
need to send the unit back to Dave Cook.

105: CO2, Site tower, Sat 16-Oct-1999 08:43:48 CDT, 40m co2 power supply taken down
The power supply unit for the 40m co2 sensor was removed
last night to run tests with the questionable co2 sensor
deployed at 20m. No data is being lost since ragwort has been
turned off.


104: ADAMS, Site tower, Sat 16-Oct-1999 08:41:09 CDT, Ragwort off the air
Ragwort was taken off the air last night because of possible
rain. The lid has to be left open because of heating. Hopefully,
we will change chassis in the next few days.

103: Hotwire, Site tower, Fri 15-Oct-1999 14:36:39 CDT, moving 40m boom
At 1337 CDT, I raised the 40m boom to mount a hotwire, #70. I lowered the
boom at 1342.

101: Sonics, Site tower, Fri 15-Oct-1999 10:37:09 CDT, Cycled power on 10m sonic on 60m tower
The 10m sonic on the 60m tower quit around 1755 local time, Oct 14.
Restarted it around 1000 local time, Oct 15.
98: Sonics, Site tower, Thu 14-Oct-1999 18:04:22 CDT, 20m sonic died; replaced w/ATI (and KH2O from Station 6)
The ATI at 20m died at 10:40 am local time, Oct 14.  We removed the
ATI and KH2O from Station 6 at 15:30 local time and installed them at 20m.  
The sonic came on line at 16:35.  
97: Hotwire, Site tower, Thu 14-Oct-1999 15:36:16 CDT, Raising 10m and 40m booms to remove hotwires
14:53:00 10m and 40m probes turned to standby in preparation for wx.

14:56:00 CDT raised the 40m boom to dismount the hotwire at 40m. Finished 
at 15:01:00.

15:20:00 raised the 10m boom to dismount that hotwire; finished at 15:24:00.

95: CO2, Site tower, Thu 14-Oct-1999 09:55:16 CDT, No licor cal gases
All of the "lo-cal" CO2 calibration gas was used after the first day that
the Licor was deployed - about 10/7?  I had identified a leak in the solenoid
valve which controlled this line in Tony's box and had tried to fix it, but
apparently was unsuccessful.  There still was plenty of "hi-cal" gas until
sometime during the last 3 days.  Yesterday morning, the "hi-cal" cylinder
also was empty.

Both cylinders are now off.  We probably should disconnect the control circuit
to let the licor sniff real air all the time, but I haven't done this yet.
I don't know how we'll generate a calibration now - perhaps we should manually
attach soda lime to get a zero from time-to-time.  I've put a call in to
Tony.
94: Hotwire, Site tower, Wed 13-Oct-1999 16:45:06 CDT, hotwire replacement/look on 55-m tower (Lundquist)
13 October 1644: Julie raisesboom at 10 m on 55-m tower to replace 10 m hotwire sensor.Julie notes there are a lot of spider webs.Installed probe 69.

1650.  10 m boom is down.

1653:  Julie reports tower vibration; feels like 0.5 Hz.

1701:  Julie is pulling up 40 m boom.Lots of spiders.

1705:  Julie installed probe 70 at 40 m.

170930:  40 m boom is down.










92: LOG, Site tower, Wed 13-Oct-1999 11:50:43 CDT, tour up 55m tower
Several people were up the 55m tower this morning:
Jielun and Shawn were working on their thercouples,
Roddy and Steve climbed up to 25m for a tour,
Julie worked on hotwires,
Gordon & Tom were replacing fiber cable.
This all between 9-11AM CDT.

89: TRH, Site tower, Tue 12-Oct-1999 18:04:34 CDT, Replaced hygrothermometer fan at 15m
Sometime this morning, perhaps 10-11 am local time, Oct 12, Steve
replaced the fan in the hygrothermometer shield at 15m.

Oct 14:

Although the temperature error was pretty small, it is now obvious
that there was a definite improvement in the 15m T following this
replacement.

Unfortunately, RH.15m is still about 2 %RH lower than RH.5m and RH.25m
during the day.  At night RH.15m is lower than both RH.25m and RH.35m.
The 15m mixing ratio is also anomolously low at all times.
85: Hotfilm, Site tower, Tue 12-Oct-1999 14:06:10 CDT,
9 Oct  1100 to 1700 Tried to find spiking problem.
9 Oct 1715  Steve/Gordon changed timing of zero circuit relative to sampling of sensors,  seems to take care of spiking problem. 
After this time the zero circuit is left on.
Before this time the zero circuit was off and base voltage was left to drift.
It was reset most days by re-installing circuit for short time. 
Spectra show spikes at 60 and 80 hz in the two cold films but not in the hot films. Spikes are small - we will probably remove them from the data after the fact.
10 Oct - 1145 to 1240 - reset timer on zero sheath covers on hotfilms.  Timer has been set wrong at zero 0300 and 1100 and 1200 hours.  Now set to cover sensors at hh:59 each hour.


82: Sonics, Site tower, Tue 12-Oct-1999 10:02:48 CDT, ati.55m power cycle
I cycled power on 55m ati which brought it back to life.  This done about
9:30 (and I was up the tower to ragwort).  This sonic has been down since
2pm yesterday - oops...
81: ADAMS, Site tower, Mon 11-Oct-1999 21:25:41 CDT, new Ironics card at ragwort

Installed new Ironics 3234 card at ragwort. pp.50m shows spikes due
to dropped characters, which I assume is due to the ironics 1624
not keeping up.



80: ADAMS, Site tower, Mon 11-Oct-1999 19:06:55 CDT, ragwort overheating
Ragwort died today at 11:34 CDT (16:34 GMT).

The LMON light on ragwort's port in the chem shelter fiber box 
was out.  Before I realized it was probably a cooling problem,
I just switched the breaker on and off.

Ragwort came up but
	siostatus yarrow
showed non-zero values in the msg ERRORS column.

I then did a mxreset and it came up with the serial card working.
It then died at 19:08 GMT.

At 23:00 I climbed the tower.  The power supply was dead - the
three red lights were out.

I redid the ventilation cardboard so that the air flows over the
power supply in addition to the 4 VME cards (but not the signal-cond
card).  Also ripped out the insulation under the top of the adam.

Finally looked at Tbox.ragwort.  A plot of Tbox.ragwort shows that
the cooler must have failed at about 06:40 CDT this morning.
Tbox went from its 10 minute cycling at about 27 degC up to 35 degC 
at 11:30 when ragwort failed.

After restarting ragwort at 18:30 Tbox shot up to 34C, so at 20:00
the lid was left propped open about 5 inches.







78: CO2, Site tower, Mon 11-Oct-1999 16:33:54 CDT, 20m H2O/CO2 reinstalled
We just put the ATDD sensor back up.  Dave Auble simply tightened connectors
inside the electronics/power supply box and removed two passive filters
added by ANL.  The data look almost believable now.

77: CO2, Site tower, Mon 11-Oct-1999 13:26:40 CDT, 20m boom up/CO2 removed
We've just removed the 20m H2O/CO2 sensor so that Dave Auble can look at
it today.  This was from about 12:30-13:00
75: Kryptons, Site tower, Mon 11-Oct-1999 11:08:28 CDT, Cleaned 5m KH2O
Cleaned 5m krypton hygrometer windows between 10:45 and 10:50 local time,
October 11.


73: Sonics, Site tower, Sun 10-Oct-1999 17:09:41 CDT, 20m ATI was at 10 samp/s
Somehow, the 20m ATI was set to 10 samp/s.  I assume that this has been the
case since set-up.  It is now at 20 samp/s.

72: Sonics, Site tower, Sun 10-Oct-1999 17:05:40 CDT, replaced 10m ATI
Something has been wrong with the 10m ATI since it was installed.  The worst
signal is just w.  I just replaced it with our "spare" which was removed from
40m earlier due to spiking and never changed.  I'm hoping that this will
work ok in the lower wind speeds at 10m.  (It ran on the bench just fine.)

Note that this entailed switching the hotwire mount, which fortunately was
straightforward.

Also note that this sonic also has a level, so the config is the same.
(Unlike when we swapped at 40m -- the sonic now there is reporting dummy
values for the levels (i.e. analog output is turned on, but there is no
signal.))

70: Sonics, Site tower, Sun 10-Oct-1999 10:49:33 CDT, 20m boom orientation
I rechecked the 20m boom this morning and found that the mount had
shifted to the south, presumably caused by the process of pushing the
boom to the north when raising it (to avoid crossed guy wires above it).

At 10:20 am local time, Oct 10, I moved the mount to its correct 
position that was marked by black electrical tape.  Unfortunately, 
the boom still appears to be angled a few degrees to the north 
relative to the other booms.
67: LOG, Site tower, Sat 09-Oct-1999 22:32:57 CDT, 5m tower down times
5m tower down:
Oct 1 most of the day
Oct 3 tower down 9 to 1100.
Releveled tower and sensors at 1000
Oct 5 1630 - 1804 tower down. Replaced 5m cold film
Oct 5 1900 - 1920  grounded cf bridges to bridge power supply



62: Tower, Site tower, Sat 09-Oct-1999 12:17:31 CDT, raised and lowered 20m boom
From 10:45 to 11:25, the boom at 20 m was raised and lowered so that we could 
switch out hot wire probes (took out #78, put in #71) at that level. 
Inspection showed that #78 broke because a tower denizen spun its web across 
the probe, probably incurring quite a jolt. 

When the boom was lowered, it appeared that the 20m boom was angling off to 
the north more than the other booms. Tightening the south guy wire did not 
change that.
60: PropVanes, Site tower, Sat 09-Oct-1999 10:58:27 CDT, swapped prop at 15m
The prop at 15m (W0002) has been reading nearly 0 speeds for quite a while.
I just spent the last hour+ swapping it with spare sensor (W0003).  I
initialized this spare in the trailer with new propeller P56159 with
a pitch of 306, reset the boom alignment (only changed by one count from
what it had initially), and set the boom angle to 97.0 deg.  (I note that
W0002 had 96.7 deg, which is off by 0.3 deg from the boom angles which
were logged.)  I put the old vane in the new sensor.  It bent slightly when
I pulled it into the tower (even when I was watching for this), but I think
I bent it back close to straight.

The speed profile appears to be better now.

59: CO2, Site tower, Fri 08-Oct-1999 17:07:37 CDT, co2.20m fixed itself?
co2.20m appears to have a reasonable signal now, despite no action being
taken on our part.  It definitely has been offscale or spiking for the
previous several days.  We'll keep watching it.

58: Sonics, Site tower, Fri 08-Oct-1999 16:50:42 CDT, Moving the 40 m boom
This entry should have been made on Tuesday, October 5. From 12:40-13:03 CDT, I raised and lowered the 40 m boom to mount a hot-wire probe next to the sonic.

57: Kryptons, Site tower, Fri 08-Oct-1999 16:21:01 CDT, new krypton at 55m
By swapping things, we found that the head of S/N 1133 is responsible for
the 50s noise (and generally bad signal) from h2o.55m.  We've swapped the
whole sensor with S/N 1389.  We'll update cal_files soon.

P.S. Obviously, we (Tom and I) were up the tower -- from about 3:00-4:10.

54: CO2, Site tower, Fri 08-Oct-1999 09:03:30 CDT, approximate licor calibration
A crude fit between the Licor and co2.40m gives:

CO2 = (Licor + 2.4)*140

in mg/m3 of CO2.  To convert to ppm:

mg/m3 * g/1000mg * mole/44g * 22.4 lit/mole * m3/10^3 lit * 10^6ppm ~ 0.5

With these factors, the cal-gas values are about:
357ppm; 330ppm; 364ppm

I will enter these values in prep.config.


52: CO2, Site tower, Fri 08-Oct-1999 08:44:28 CDT, Licor co2 installed
Yesterday, I found that the Licor problem was wrong pins connected to the
CO2 analog output channel and corrected this problem about noon.  Signal + 
was connected to power ground (pin7) and Signal return to open (pin4).  
The signal now appears to be working.  At that time, I also opened the cal
gas cylinders and installed the valve control lines.  I see 2 calibrations
in the data at about 4PM and 4AM last night.


51: ADAMS, Site tower, Thu 07-Oct-1999 17:13:10 CDT, daisy problems
Daisy is sick.

Configuration:

  daisy
  chicory
  datel
  chassis #4

Did lots of switching cards around.  

Indications also pointed to cooling problems, because the adam would 
work fine and then die 1 half hour after closing the lid.
(Daisy's lid is missing insulation, which should keep it cooler.)

The new matrix card has its own problems.  After a while it would
choke on a bad sample (chan==0,length==0).  This appears to be
some sort of memory corruption problem.

The new ironics cards may work now, but perhaps run hot.
They have 4 meg of RAM (the others have 1 meg).

Ventilation was improved by duct taping cardboard in the fan "horn"
with gaps only at the ends so that the cool air is directed only
to the 4 cards on the left (matrix,ironics,serial breakout and datel)
and to the power supply on the right.

Here is what is working:

  daisy matrix (S/N #9010
  chicory ironics (#014029)
  daisy's original datel (#440690)
  PSU power supply
  chassis #4


48: CO2, Site tower, Tue 05-Oct-1999 19:41:05 CDT, co2 turned off
turned off Licor (includeing pump and cal gases), since it has been reporting
bad values.  This done about 6pm local.  It has been running bad for the last
day and a half.

44: Sonics, Site tower, Mon 04-Oct-1999 10:53:32 CDT, Movement of booms
From 0930 to 1045, Horst and Lundquist raised and lowered the beams at 
40m (to remove a hotwire probe shield and support), 20 m (to mount a probe),
and 10 m (to switch out shield/support and mount a probe), in that order.

39: ADAMS, Site tower, Thu 30-Sep-1999 12:30:00 CDT, ADAM status
Swapped the matrix board in ragwort with daisy's.

Results:

  daisy could not talk to datel (HSTSTA errors)

  ragwort is sick, will not stay up. 
  After a while it does not see ironics board and mxreset does not
  cure it. A power cycle will work for a while

Then on 10/1
  put the daisy matrix board back in daisy:  works

  put the PSU matrix board into ragwort.  The EPROM is configured
  for 1 meg RAM without PPP.
  
  The PSU ragwort booted, but then "siostatus yarrow" showed msg errors.
  After a while it died.  Same symptoms as the ragwort matrix.


   

38: ADAMS, Site tower, Wed 29-Sep-1999 21:07:06 CDT, ADAM status

Power to tower was down last night due to a rabbit's chewing.   

I was distracted so didn't monitor adams much today. After power
was restored the adams were up except for ragwort.

Ragwort was dead (status light orange, Ironics halted). 

Replaced the Ironics 1624 in ragwort (yarrow) with a new iv3234.
32: ADAMS, Site tower, Tue 28-Sep-1999 09:15:38 CDT, ADAM status
Lots of rain last night with lightning.  Must have had power outages
because the adams all rebooted at the same times.  Toucan reboots
cause breaks in the PAM archive.

The log from the UPS in the trailer showed some entries for these
times also.

01:01:50	cosmos  ?? (probably not a power outage)
01:55:00	cosmos,daisy,marigold,ragwort,toucan
02:20:30	cosmsos,daisy,marigold, toucan
02:21:30	cosmos,daisy,marigold,ragwort
03:10:50	cosmos,daisy,marigold,ragwort,toucan
12:01:30	cosmos,daisy,marigold,ragwort,toucan

Everything was running in the morning, except the serial channels
on ragwort.  "siostatus yarrow" showed errors in the "msgs" column.
mxreset would not bring the serial channels back.  Power cycling
ragwort from the base of the tower did the trick.

The last network fix to ragwort was swapping the fiber transceivers
with cosmos.  That must have helped because we haven't had any
unexplained reboots since then. 











29: ADAMS, Site tower, Sun 26-Sep-1999 08:54:50 CDT, ADAM status
cosmos:
  ran overnight
daisy: 
  ran overnight
marigold:
  ran overnight
ragwort:
  dead this morning.
  From archive:
     died at 06:18 GMT (01:18 am CDT)
     restarted itself at 06:25 GMT.
     died at 08:30 GMT
     restarted itself at 11:43 died immediately
     rebooted at 14:43 (9:43) by human

  Log file on toshiba was empty - serial
  connector fell off after setting it up.

  Swapped fiber transceiver with the one from cosmos.  Rebooted at
  15:47 GMT

  








28: Sonics, Site tower, Sat 25-Sep-1999 16:43:03 CDT, sonic parameters
adam	chan	height	type	owner	serial

cosmos	200	2.5m	CSAT	Miller	
cosmos	201	5m	CSAT	Miller
cosmos	204	10m	ATI	NCAR	980505 (w/levels)
cosmos	205	20m	ATI	NCAR	980202
marigold 200	30m	CSAT	Mahrt
ragwort	200	40m	ATI	NCAR	980504 (w/levels)
ragwort	201	50m	CSAT	Mahrt
ragwort	202	55m	ATI	PNL	980415

All ATIs:
Sample Options Menu\r\n
A. Median Filter (Off)\r\n
B. Data Quality Algorithm (Off)\r\n
C. Shadow Correction Algorithm (Off)\r\n
D. Remove RH from Temp Calculation (Off)\r\n

27: ADAMS, Site tower, Sat 25-Sep-1999 16:20:53 CDT, ADAM status, turned off chem shelter AC
cosmos:
  OK
daisy: 
  crashed and auto rebooted again.
  Switched to other two fibers in cable, from green/blue to orange/brown.
marigold:
  running fine and dandy.

  Then at 7:10 pm CST (00:11 GMT) it went nuts, flooding the net 
  with traffic (removing its fiber from the hub quieted things down).
  Did not look at its console output.  A power cycle cured things.
ragwort:
  crashed and auto rebooted once or twice. When visited on tower,
  console was not responding, and status light was dull orange.
  Cable to Toshiba was loose, so no info was available.

  Rebooted ragwort with toshiba attached to console. After a while
  saw "med0: excessive collision" errors.

  Switched to other two fibers in cable, from orange/brown to green/blue.


Also, turned off the chem shelter air conditioner, in case the
problem could be power related.
The AC is on the same transformer as ragwort,daisy,marigold.







24: ADAMS, Site tower, Sat 25-Sep-1999 12:21:34 CDT, ADAM status
cosmos:
  Fixed its problems this morning.  Software bug in krypton calibration
  function couldn't handle empty lines in the cal file and would
  infinitely loop, hanging whatever host was running covar, cockpit, or
  prep.  Cosmos is now ingesting all its channels.
daisy: 
  rebooted itself this morning for no apparent reason.  
marigold:
  running fine and dandy.
ragwort:
  On Friday, 9/24, replaced the 40m fiber borrowed from ISS with
  a 4-fiber 300m cable.  The 40m cable is still tie-wrapped to the tower.
  This was done to fix problem losing fiber link (LMON light on Dyna FL/4
  port would go out and ragwort would reboot).

  ragwort died this morning.   The halt light on the Ironics card was red.
  The light on the Matrix was dim red I believe.  Power glitch?
  Rebooted, with the toshiba logging the console.


22: Tower, Site tower, Fri 24-Sep-1999 09:06:49 MDT, 55m tower guy wire tensions

    55 meter tower guy wire tensions:


   September 13, 1999  (just after retensioning)

Guy level    Nominal     NW corner     SW corner     SE corner     NE corner
             Tension
              (lbs.)

   18'     425-475        475           475           575           450

   36'     325-375        260           160           300           350

   54'     225-275        200           250           250           215

   78'     575-625        690           510           675           625

   102'    525-575        600           525           700           600

   126'    625-675        650           675           625           625

   150'    575-625        650           510           525           700

   174'    725-775        750           725           775           650

   186'    675-725        590           675           725           675


   September 22, 1999  (weekly recheck)

                          NW            SW            SE            NE

   18'     425-475        450           450           425           370

   36'     325-375        265           275           310           280

   54'     225-275        180           190           250           200 

   78'     575-625        650           475           600           525

   102'    525-575        550           515           650           540

   126'    625-675        570           650           540           600

   150'    575-625        625           530           500           670

   174'    725-775        700           700           700           650

   186'    675-725        520           685           735           625


differences in guy tensions at the same level and at different times could
be caused by many different factors - wind loading, slipage of guy clamps,
slipage of anchors due to wind or moisture, etc.


 September 29, 1999 weekly recheck

                          NW            SW            SE            NE

   18'     425-475       375           400           390           290
   36'     325-375       260           250           300           300
   54'     225-275       200           180           230           210
   78'     575-625       600           440           500           500
   102'    525-575       510           540           525           615
   126'    625-675       570           600           520           550
   150'    575-625       580           500           495           625
   174'    725-775       725           680           670           600
   186'    675-725       550           600           630           625

 October 8, 1999 weekly recheck by Marcel V.
			NW		SW		SE		NE
18'	425-475		490		425		400		300
36'	325-375		250		250		300		450
54'	225-275		280		175		250		200
78'	575-625		575		475		540		460
102'	525-575		525		510		500		590
126'	625-675		550		610		525		550
150'	575-625		600		500		490		625
174'	725-775		700		675		700		600
186'	675-725		500		690		700		610


 October 18, 1999 weekly check by Steve S. (unit 2150)
			NW		SW		SE		NE
18'	425-475		410		400		425		290
36'	325-375		240		275		325		300
54'	225-275		200		175		240		200
78'	575-625		600		500		550		500
102'	525-575		550		575		550		575
126'	625-675		550		650		575		550
150'	575-625		575		525		500		650
174'	725-775		700		700		725		625
186'	675-725		500		700		675		600

 October 26, weekly check by Tony D and Sean  ? Wind from S (Unit 2150)
                        NW              SW              SE              NE
18'     425-475         340             380             400             280
36'     325-375         300             270             300             330
54'     225-275         200             180             260             230
78'     575-625         540             460             520             470
102'    525-575         480             520             510             560
126'    625-675         500             600             560             510
150'    575-625         550             520             500             610
174'    725-775         660             660             710             590
186'    625-725         510             710             660             600